Grade 13Mechanics

A body of mass 400 g fall and strickes the ground with velocity of 30ms-1.how much is the kinetics energy of the bodyat the time it strikes the ground

Afaq ahmad
5 Years agoGrade 13

2 Answers

abhay gupta
5 Years ago
Dear student you can find you answer here,

We need more info like height magnitude to solve this question

after getting value of height
here’s the solution

Since WORK DONE = GAIN IN KINETIC ENERGY
=> W = (K.E)final– (K.E)initial
=> mgh =(K.E)final – ½ x 0.4 x 302
=> 0.4 x 9.8 x h =(K.E)final – 180
=>(K.E)final = 39.2h + 180

Askiitians expert
Ankit
Vibekjyoti Sahoo
5 Years ago

KE = 1/2mv²; where m is mass in kg, and v is velocity in m/s.

Convert mass to kg.

400 g × 1 kg/1000 g = 0.4 kg

KE = 1/2 × 0.4 kg × (30 m/s)² = 180 kg•m²/s² = 180 Joules